# Low-Cost FlashAttention with Fused Exponential and Multiplication Hardware Operators

### v1

@@ I Introduction
  • The discrete steps of exponential function evaluation and the subsequent vector floatint-point multiplication inherent in FlashAttention kernel are fused to one simplified ExpMul hardware operator.
@@ II-A Attention Mechanism
8:         $\vec{o}_{i}\leftarrow\vec{o}_{i-1}+e^{s_{i}-m_{N}}\cdot\vec{v}_{i}$
9:         $\ell_{i}\leftarrow\ell_{i-1}+e^{s_{i}-m_{N}}$
@@
In the following, the output vector $\vec{o}_{i}$ is computed incrementally as the weighted sum of each value vector and the corresponding exponential score decremented by the maximum score. In parallel, the sum of all exponentials $\ell_{i}$ is accumulated. The final attention vector for one query vector $q$ is computed by dividing the final output vector by the corresponding sum-of-exponents $\ell_{N}$.
@@ IV-B Fused Exponential-Vector Multiplication Operator
To simplify the computation, we convert the value of $X$ to its fixed-point equivalent $\hat{X}$. In general, converting a floating-point number to its fixed-point representation requires many integer bits to accommodate the wide range of floating-point arithmetic [28]. However, in this case, we can safely clip $X$ to a much smaller dynamic range, as it will only be used for the computation of $e^{X}$ for negative values of $X$. In this range, $e^{X}$ quickly converges to 0, even for small negative $X$ values, i.e., $e^{-15}=3.1\cdot 10^{-7}$. In other words we can convert $X$ to a fixed point number with a small number of bits without the need to consider its full dynamic range. Thus, instead of fully quantizing $X$ we simply constrain its value before logarithmic quantization.
@@
Converting to a fixed point number the clipped value of $X$ requires a small number of bits. For this conversion, we use a 16-bit fixed point number to keep the corresponding hardware overhead small. Since $X$ is going to be multiplied by a constant value its range after clipping will change to [21.64, 0]21.64 0[-21.64,\ 0][ - 21.64 , 0 ]. To account for this new range we assign 6 bits for the integer part and the rest for the fraction part.
@@
Since $\hat{X}$ is a fixed point number the multiplication with the constant value $\log_{2}e$ can be approximated by integer shift–and–add operations [26, 25]. Following the approach presented in [26], we get that:
@@
Multiplying the floating point number $V=(-1)^{S_{V}}2^{E_{V}-b}(1+M_{V})$, where $S_{V},E_{V},M_{V}$ represent its sign, exponent and mantissa field, respectively, with $2^{-\hat{L}}$ in (8) is equivalent to subtracting $\hat{L}$ from the exponent of $V$. In case of an overflow the resulting FP number is set to 0. More specifically,
@@
Figure 2: The FlashAttention-2 kernel optimized with the proposed ExpMul operators. According to Alg. 3 each ExpMul operator involves clipping, integer shift-and-add operations and an exponent increment producing directly floating point results without further dequantization.
@@
The parallel hardware organization of FlashAttention-2 kernel employing the proposed ExpMul operators is shown in Fig. 2. The organization of the new hardware unit is the same as the original FlashAttention-2 accelerator shown in Fig. 1. However, we can highlight one key difference. The exponent function hardware operators and floating point multipliers are replaced by a single ExpMul unit which performs those two operations in a fused manner using integer only arithmetic and preparing the result in floating point as needed by subsequent operations.
@@ V-A Verification of LLM accuracy
Since ExpMul involves logarithmic quantization, it introduces a numerical approximation. To verify that this approximation does not hinder the capabilities of a Large Language Model (LLM), we run ten representative NLP benchmarks from the GLUE [16] dataset, using Google’s Flan-T5 LLM. For comparison, we considered four implementations: (a) Models ’FP32’ and ’BF16’ use the default attention mechanism of T5 and compute all operations including separate multiplications and function evaluation using single-precision and BFloat16 [29] floating point arithmetic; (b) Models ’FP32-ExpMul’ and ’BF16-ExpMul’ employ floating point arithmetic for all parts of the model except for the proposed ExpMul operator, where computations are done in 16-bit integer arithmetic, as discussed in Section IV-B. To run inference with the LLM we utilized Microsoft’s PromptBench workflow [30].
@@ V-B Hardware Savings
To evaluate the hardware savings of the proposed ExpMul hardware operators we implemented two variants of the FlashAttention-2 kernel. The two designs correspond to the main block shown in the foreground of Figs. 1 and Fig. 2, respectively. The implementation involves various hidden dimension sizes ($d$) and two floating-point formats: FP32 and BFloat16. The baseline FlashAttention kernel performs exponent and multiplication operations separately. Exponent function evaluation is implemented with piece-wise linear approximation in the reduced input range inherent to attention and used also for the ExpMul operators.
@@
As shown in Fig. 3, utilizing ExpMul operators reduces the hardware area by more than 28.8% on average in all examined cases. These saving come from the simplified computation kernel that does not require area expensive floating point multipliers and or additional exponential function evaluations. In the output update module, $d$ multiplication and exponentials are converted to ExpMul operators, while reforming the floating-point output requires only $d$ exponent subtractions.
Simplifying hardware architecture improves energy efficiency, reducing power consumption by over 17.8% on average. As the hidden dimension $d$ increases, power savings increase due to replacing power-hungry floating point multipliers with simpler ExpMul operators that use integer arithmetic.

### v2

@@ I Introduction
  • The discrete steps of exponential function evaluation and the subsequent vector floating-point multiplication inherent in FlashAttention kernel are fused to one simplified ExpMul hardware operator.
@@ II-A Attention Mechanism
8:         $\ell_{i}\leftarrow\ell_{i-1}+e^{s_{i}-m_{N}}$
9:         $\vec{o}_{i}\leftarrow\vec{o}_{i-1}+e^{s_{i}-m_{N}}\cdot\vec{v}_{i}$
@@
In the following, the output vector $\vec{o}_{i}$ is computed incrementally as the weighted sum of each value vector and the corresponding exponential score decremented by the maximum score. In parallel, the sum of all exponentials $\ell_{i}$ is accumulated. The final attention vector for one query vector $\vec{q}$ is computed by dividing the final output vector by the corresponding sum-of-exponents $\ell_{N}$.
@@ IV-B Fused Exponential-Vector Multiplication Operator
To simplify the computation, we convert the value of $X$ to its fixed-point equivalent $\hat{X}$. In general, converting a floating-point number to its fixed-point representation requires many integer bits to accommodate the wide range of floating-point arithmetic [28]. However, in this case, we can safely clip $X$ to a much smaller dynamic range, as it will only be used for the computation of $e^{X}$ for negative values of $X$. In this range, $e^{X}$ quickly converges to 0, even for small negative $X$ values, i.e., $e^{-15}=3.1\cdot 10^{-7}$. In other words we can convert $X$ to a fixed-point number with a small number of bits without the need to consider its full dynamic range. Thus, instead of fully quantizing $X$ we simply constrain its value before logarithmic quantization.
@@
Converting to a fixed-point number the clipped value of $X$ requires a small number of bits. For this conversion, we use a 16-bit fixed-point number to keep the corresponding hardware overhead small. Since $X$ is going to be multiplied by a constant value its range after clipping will change to [21.64, 0]21.64 0[-21.64,\ 0][ - 21.64 , 0 ]. To account for this new range we assign 6 bits for the integer part and the rest for the fraction part.
@@
Since $\hat{X}$ is a fixed-point number the multiplication with the constant value $\log_{2}e$ can be approximated by integer shift–and–add operations [26, 25]. Following the approach presented in [26], we get that:
@@
Multiplying the floating-point number $V=(-1)^{S_{V}}2^{E_{V}-b}(1+M_{V})$, where $S_{V},E_{V},M_{V}$ represent its sign, exponent and mantissa field, respectively, with $2^{-\hat{L}}$ in (8) is equivalent to subtracting $\hat{L}$ from the exponent of $V$. In case of an overflow the resulting FP number is set to 0. More specifically,
@@
Figure 2: The FlashAttention-2 kernel optimized with the proposed ExpMul operators. According to Alg. 3 each ExpMul operator involves clipping, integer shift-and-add operations and an exponent increment producing directly floating-point results without further dequantization.
@@
The parallel hardware organization of FlashAttention-2 kernel employing the proposed ExpMul operators is shown in Fig. 2. The organization of the new hardware unit is the same as the original FlashAttention-2 accelerator shown in Fig. 1. However, we can highlight one key difference. The exponent function hardware operators and floating-point multipliers are replaced by a single ExpMul unit which performs those two operations in a fused manner using integer only arithmetic and preparing the result in floating-point as needed by subsequent operations.
@@ V-A Verification of LLM accuracy
Since ExpMul involves logarithmic quantization, it introduces a numerical approximation. To verify that this approximation does not hinder the capabilities of a Large Language Model (LLM), we run ten representative NLP benchmarks from the GLUE [16] dataset, using Google’s Flan-T5 LLM. For comparison, we considered four implementations: (a) Models ’FP32’ and ’BF16’ use the default attention mechanism of T5 and compute all operations including separate multiplications and function evaluation using single-precision and BFloat16 [29] floating-point arithmetic; (b) Models ’FP32-ExpMul’ and ’BF16-ExpMul’ employ floating-point arithmetic for all parts of the model except for the proposed ExpMul operator, where computations are done in 16-bit integer arithmetic, as discussed in Section IV-B. To run inference with the LLM we utilized Microsoft’s PromptBench workflow [30].
@@ V-B Hardware Savings
To evaluate the hardware savings of the proposed ExpMul hardware operators we implemented two variants of the FlashAttention-2 kernel. The two designs correspond to the main block shown in the foreground of Fig. 1 and Fig. 2, respectively. The implementation involves various hidden dimension sizes ($d$) and two floating-point formats: FP32 and BFloat16. The baseline FlashAttention kernel performs exponent and multiplication operations separately. Exponent function evaluation is implemented with piece-wise linear approximation in the reduced input range inherent to attention and used also for the ExpMul operators.
@@
As shown in Fig. 3, utilizing ExpMul operators reduces the hardware area by more than 28.8% on average in all examined cases. These savings come from the simplified computation kernel that does not require area expensive floating-point multipliers and or additional exponential function evaluations. In the output update module, $d$ multiplication and exponentials are converted to ExpMul operators, while reforming the floating-point output requires only $d$ exponent subtractions.
Simplifying hardware architecture improves energy efficiency, reducing power consumption by over 17.8% on average. As the hidden dimension $d$ increases, power savings increase due to replacing power-hungry floating-point multipliers with simpler ExpMul operators that use integer arithmetic.